# Realised Volatility Forecasting: Machine Learning via Financial Word Embedding

### v5

@@ 1 Introduction
Most studies on realised volatility forecasting use historical realised volatility as the primary source of data for prediction within a linear framework. Heterogeneous autoregressive (HAR) models (Corsi 2009) are well-known, straightforward linear models for realised volatility forecasting that parsimoniously capture volatility persistence across multiple horizons. Further developments expanded the HAR-family with variations like HAR-J (HAR with jumps) and CHAR (continuous HAR) from Andersen et al. (2007b), SHAR (semivariance-HAR) from Patton and Sheppard (2015), and the HARQ model from Bollerslev et al. (2016). Bollerslev et al. (2018) extend these models by demonstrating that exploiting cross-asset volatility similarities improves out-of-sample forecasting performance, while Patton and Zhang (2025) show that further gains can be achieved by customising realised volatility measures. Also, during the last decade, there has been a growing number of studies on the theory and application of machine learning (ML) in finance. Recent evidence shows that ML models can outperform traditional financial models in asset pricing and return prediction (Gu et al. 2020), short-term price forecasting using high-frequency data (Sirignano and Cont 2019), and text-based financial forecasting using news articles (Adämmer and Schüssler 2020). More recently, Bybee et al. (2024) employ a latent Dirichlet allocation (LDA) topic model to analyse the textual content of Wall Street Journal articles, demonstrating that business news attention closely tracks economic activity and improves forecasts of macroeconomic dynamics.
@@
In the context of realised volatility forecasting, Sermpinis et al. (2013), Rahimikia and Poon (2020), Christensen et al. (2023), and Han et al. (2025) show that ML models can outperform traditional econometric approaches in predicting volatility. Moreno-Pino and Zohren (2024) introduce a realised volatility forecasting model using dilated causal convolutions and find that it improved prediction accuracy compared to traditional methods. Li and Tang (2025) introduce an automated volatility forecasting system that leverages a broad array of features and multiple algorithms, achieving superior out-of-sample forecasting performance across different horizons. A closely related strand of research shows that textual news can be processed automatically to quantify market reactions, with Groß-Klußmann and Hautsch (2011) demonstrating that machine-read news contains economically meaningful information for high-frequency market activity and volatility. Notwithstanding these advances, other studies caution that the empirical performance of ML models can be unstable and context-dependent (Hillebrand and Medeiros 2010, Audrino and Knaus 2016, Branco et al. 2024, Audrino and Chassot 2025). Also, beyond realised volatility, there are a handful of studies that specifically focus on NLP and large language models (LLMs), such as Van Binsbergen et al. (2024), who employ ML techniques using word embeddings to construct a 170-year-long measure of economic sentiment from 200 million pages of US local newspapers, demonstrating its predictive power for GDP growth, employment, and monetary policy decisions. More recently, studies such as Chen et al. (2022), Huang et al. (2023), and Rahimikia and Drinkall (2024) have incorporated information extracted from advanced LLMs for various financial tasks.
@@
Taken together, prior work has advanced realised volatility forecasting primarily by refining models that exploit the persistence of volatility and by expanding the predictor set with additional realised measures or broader ML feature sets. At the same time, text-as-data research shows that NLP and LLMs can extract economically meaningful signals from large corpora, but their role as dedicated, standalone forecasting models for realised volatility remains underexplored. This motivates four objectives that are central for both researchers and practitioners. First, we assess whether news alone, without lagged realised volatility or other market variables, can generate competitive realised volatility forecasts, and whether general-purpose language representations are sufficiently competitive for this task or specialised, finance-oriented NLP components are required. In doing so, we emphasise word embeddings as the core representation that converts unstructured news into quantitative inputs by encoding semantic and contextual similarity in a low-dimensional vector space; such embeddings also underpin modern LLMs as their representation layer, making them a transparent and practically relevant starting point for state-of-the-art NLP models in forecasting. Second, we examine whether any forecasting gains from news are stable across time and market conditions or concentrated in particular regimes. Third, we identify which components of news content drive predictability, distinguishing whether improvements arise primarily from stock-related or general news. Fourth, we evaluate the incremental value of combining news-based signals with standard volatility-history benchmarks relative to using either information set in isolation. We address these objectives by developing a modern yet operationally simple NLP model that maps news directly into volatility forecasts. We evaluate it both as a news-only forecaster and as a complement to standard RV models, and use explainability analysis to identify the phrases and themes most responsible for predictive performance. We also extend the analysis to assess the economic gains of the proposed models.
@@ 3 Financial Word Embedding
Although general-purpose word embeddings have demonstrated strong performance in broad language tasks, it is important to consider the potential benefits of developing specialised word embeddings and to conduct a comparative evaluation with general-purpose alternatives at the language level. Financial language exhibits distinctive usage patterns, institutional references, and contextual relationships that are often absent from general text corpora. As a result, many tokens may carry meanings, associations, or economic implications that differ substantially from their usage in non-financial contexts. Using general-purpose word embeddings without prior validation may therefore introduce measurement error into text-based variables and weaken their economic interpretability. This consideration motivates the development and evaluation of specialised word embeddings tailored to financial language before employing them in downstream financial applications. Section 3.1 describes the steps from data collection to the training of the financial word embedding, together with the associated design considerations, and Section 3.2 subsequently presents the evaluation of both general-purpose and financial word embeddings using representation analysis and multiple language level benchmarks.
@@ 3.1 Design
Moving to corpus size, the Google Word2Vec, WikiNews, and FinText corpora contain approximately 100 billion, 16 billion, and 4.32 billion words, respectively, implying that FinText relies on a substantially smaller corpus than these general-purpose word embeddings. Nevertheless, the FinText dataset expands substantially over time, with the number of monthly samples and the total word count exhibiting clear long-run growth over the 15-year period covered by the news archive, albeit with noticeable time variation, as illustrated in Figure 1. Following all pre-processing steps, the resulting FinText corpus comprises 2,733,035 unique tokens.
@@
3.2 Language Level Evaluation
@@
Table 1 compares four versions of FinText against general-purpose word embeddings based on the Google Analogy benchmark. The Google Analogy benchmark is a widely used dataset for evaluating the quality of word embeddings. Each section in the Google Analogy benchmark contains a set of analogies. For example, under the ‘capital-common-countries’ section, the word embedding is challenged with questions such as ‘London to England is like Paris to ?’. From Table 1, it is apparent that, except for ‘currency’ and ‘gram1-adjective-to-adverb’, WikiNews achieves the highest predictive accuracy. The overall evaluation score supports this observation. For this general-purpose benchmark, FinText is outperformed by Google under Word2Vec and by WikiNews under FastText. The individual and overall scores for FinText indicate that the skip-gram model performs better than CBOW.
@@ 4 Realised Volatility Forecasting
A substantial body of literature establishes that financial market volatility is fundamentally driven by the arrival and interpretation of information. Seminal contributions such as Engle and Ng (1993) and Andersen and Bollerslev (1998), together with subsequent intraday evidence, show that information arrival and return innovations are closely linked to volatility clustering and sudden volatility episodes (Gallo and Pacini 2000, Kalev et al. 2004). More broadly, this line of research emphasises that volatility dynamics reflect heterogeneous information flows operating across different horizons, which in turn generate the persistent behaviour commonly observed in financial markets (Andersen et al. 2007b). However, most empirical volatility forecasting models incorporate this information only indirectly, relying on past volatility measures as reduced-form summaries of how markets have responded to information arrivals. Although this approach has proven empirically successful, it abstracts from the underlying informational content and treats volatility persistence primarily as a time-series phenomenon. This provides motivation for testing the use of textual news as a direct proxy for information arrival.
@@ 5.2 Forecasting Performance of Ensemble Models
Although Section 5.1 shows that NLP models deliver promising results and that news content contains meaningful information for predicting RV across different market regimes, these models underperform the HAR-family of models in terms of standard forecasting metrics, particularly with respect to the magnitude of performance improvements. Combining forecasts from multiple models provides diversification gains and can improve predictive performance when individual forecasts rely on heterogeneous information sets or are subject to model uncertainty (Bates and Granger 1969, Timmermann 2006). This naturally raises the question of whether, rather than serving as a replacement for the HAR-family of models, news-based NLP forecasts can provide complementary information. To this end, we define a straightforward ensemble model that combines forecasts from the CHAR model, which is the best-performing specification within the HAR-family in Section 5, with forecasts generated by the NLP models. For each day, we compute the average of the two forecasts. This approach integrates the persistent dynamics captured by historical RV with short term information conveyed by recent news, potentially enhancing the robustness of daily RV forecasts.
@@
Taken together, the ensemble results clarify and strengthen the conclusions drawn from the standalone NLP analysis in Section 5.1. While the NLP models typically underperformed the best HAR-family specifications in terms of loss ratios, the ensemble approach successfully converts the informational content embedded in news text into statistically meaningful forecasting gains. This complementarity is strongest for stock-related news, particularly in high volatility days, consistent with the earlier finding that stock-related news is most informative during periods of market stress. For general hot news, the ensemble results mirror the regime dependence identified previously, with benefits concentrated on normal volatility days. Across both news types, the ensemble evidence reinforces the earlier conclusion that performance improvements are driven primarily by the combination of heterogeneous information sources rather than by increased model complexity. More importantly, these results indicate that news-based signals provide incremental predictive information for RV rather than fully substituting for volatility-history dynamics. Finally, similar to standalone NLP models in Section 5.1, specialised word embeddings perform better for stock-related news, while general-purpose word embeddings perform better for general hot news.
@@ 5.4 Explainable AI (XAI)
In LABEL:RV_movers_stock_related, we grouped the stock-related volatility drivers into ‘Analyst opinion’, ‘Event’, ‘Verb’, ‘Market’, ‘Abbreviation’, ‘Country/Company’, ‘Announcement’, ‘Numeric’, ‘Calendar’, ‘Insider’, and ‘Mixed’. [^27] It is important to note that this classification is based on our judgement. However, we believe that reassigning some of the n-grams between groups would not alter the fundamental informational content of these groups. The key findings can be summarised as follows: (i) ‘Analyst opinion’ and ‘Event’ contain the majority of volatility driver n-grams. This outcome is anticipated, as it clearly demonstrates the importance of analyst opinions concerning a company’s earnings calls and financial reports. Among others, popular n-grams in this group include registers, announces, files, raises, and surrenders. (ii) ‘Market’ includes market-related n-grams such as stocks to buy, premarket, and stock market opens. (iii) China is the only country in the volatility drivers list for stock-related news, underscoring the relevance of news related to this country. The remaining classes contain fewer n-grams and exhibit less commonality; however, it is evident that these n-grams in stock-related news convey varying levels of information about stocks, the market, and the economy.
@@
LABEL:RV_movers_general_hot presents the volatility drivers in general hot news, grouped into ‘Person’, ‘Place’, ‘Legal entity’, ‘Level’, ‘Verb’, ‘Index’, ‘Data’, ‘Numeric’, and ‘Mixed’, and is revealing in several ways. The ‘Person’ category is dominated by prominent political and policy figures, including U.S. presidents Donald Trump, Barack Obama, and Joe Biden, New York City mayor Bill de Blasio, senior political actors such as Mark Meadows, Mitch McConnell, Mike Pompeo, and Nancy Pelosi, as well as key Federal Reserve officials, including Chair Jerome Powell, Janet Yellen, and Federal Reserve Bank presidents James Bullard of St. Louis, Loretta Mester of Cleveland, John C. Williams of New York, Neel Kashkari of Minneapolis, and Raphael Bostic of Atlanta, underscoring the central role of political and monetary policy communication in driving market volatility. International political figures such as Dominic Cummings in the United Kingdom and Kim Jong-un of North Korea also appear as important volatility drivers, while the presence of Jim Cramer, a prominent financial media commentator, is notable but unsurprising given the influence of news coverage on market sentiment. In contrast, the appearance of William G. Kaelin, a Nobel prize winning physician scientist, and Reinhard Genzel, an astrophysicist, does not align with the economic and financial focus of the remaining n-grams in this category, illustrating that XAI methods are not error free and that their outputs must be interpreted with caution.
@@
Further analysis of LABEL:RV_movers_general_hot highlights the significance of places in our analysis and underscores the role of a group of countries, including China, as drivers of volatility. The term ‘Legal entity’ encompasses a variety of major offices, departments, commissions, and companies. The presence of the health organization and CDC (centres for disease control and prevention) could likely be attributed to the COVID-19 pandemic. [^28] As expected, the word COVID-19 does not appear, as it was not in the corpus before 2015. The next significant class, with a high number of n-grams, is labelled ‘Level’, encompassing a range of levels and changes, percentages, currency values, and specific terms like below, above, fall, and under, all referencing certain quantitative expectations. Moving to the subsequent groups, similar to stock-related news, the ‘Verb’ and ‘Numeric’ groups in general hot news underscore the relevance of these particular verbs and numbers as volatility drivers. The ‘Mixed’ group includes a variety of n-grams, such as SPAC [^29] Special-purpose acquisition company., Payroll-tax cut, Airstrike, Shutdown, Coalitions, Crisis, Trade Speech, Attorney General, and Hearing. Finally, the ‘Index’ and ‘Data’ groups underscore the importance of changes in the S&P500 index and various financial and economic indicators such as inflation, GDP, and deficit as key n-grams. A detailed analysis of all n-grams is beyond the scope of this study; however, we believe that most of these n-grams align with expectations.

### v6

@@ 1 Introduction
Most studies on realised volatility forecasting use historical realised volatility as the primary source of data for prediction within a linear framework. Heterogeneous autoregressive (HAR) models (Corsi 2009) are well-known, straightforward linear models for realised volatility forecasting that parsimoniously capture volatility persistence across multiple horizons. Further developments expanded the HAR-family with variations such as HAR-J (HAR with jumps) and CHAR (continuous HAR) from Andersen et al. (2007b), SHAR (semivariance-HAR) from Patton and Sheppard (2015), and the HARQ model from Bollerslev et al. (2016). Bollerslev et al. (2018) extend these models by demonstrating that exploiting cross-asset volatility similarities improves out-of-sample forecasting performance, while Patton and Zhang (2025) show that further gains can be achieved by customising realised volatility measures. An additional strand of the literature emphasises measurement error in realised volatility and its implications for modelling and forecasting. For instance, Cipollini et al. (2021) show that such errors are heteroskedastic and can bias standard models, proposing robust specifications to address this. Moreover, Gallo and Otranto (2015) show that modelling time-varying, regime-dependent average volatility improves forecast performance. Over the last decade, research on the theory and application of machine learning (ML) in finance has grown substantially. Recent evidence shows that ML models can outperform traditional financial models in asset pricing and return prediction (Gu et al. 2020), short-term price forecasting using high-frequency data (Sirignano and Cont 2019), and text-based financial forecasting using news articles (Adämmer and Schüssler 2020). More recently, Bybee et al. (2024) employ a latent Dirichlet allocation (LDA) topic model to analyse the textual content of Wall Street Journal articles, demonstrating that business news attention closely tracks economic activity and improves forecasts of macroeconomic dynamics.
@@
In the context of realised volatility forecasting, Rahimikia and Poon (2020) and Christensen et al. (2023) show that ML models can outperform traditional econometric approaches in predicting volatility. Ma et al. (2023) show that combining macroeconomic and financial predictors with variable selection significantly improves volatility forecasting performance. Moreno-Pino and Zohren (2024) introduce a realised volatility forecasting model using dilated causal convolutions and find that it improves prediction accuracy compared to traditional methods. Li and Tang (2025) introduce an automated volatility forecasting system that leverages a broad array of features and multiple algorithms, achieving superior out-of-sample forecasting performance across different horizons. A closely related strand of research shows that textual news can be processed automatically to quantify market reactions, with Groß-Klußmann and Hautsch (2011) demonstrating that machine-read news contains economically meaningful information for high-frequency market activity and volatility. Notwithstanding these advances, other studies caution that the empirical performance of ML models can be unstable and context-dependent (Hillebrand and Medeiros 2010, Audrino and Knaus 2016, Branco et al. 2024, Audrino and Chassot 2025). Beyond realised volatility, there are a handful of studies that specifically focus on NLP and large language models (LLMs), such as Van Binsbergen et al. (2024), who employ ML techniques using word embeddings to construct a 170-year-long measure of economic sentiment from 200 million pages of US local newspapers, demonstrating its predictive power for GDP growth, employment, and monetary policy decisions. More recently, studies such as Chen et al. (2022), Huang et al. (2023), and Rahimikia and Drinkall (2024) have incorporated information extracted from advanced LLMs for various financial tasks.
@@
Taken together, prior work has advanced realised volatility forecasting primarily by refining models that exploit the persistence of volatility and by expanding the predictor set with additional realised measures or broader ML feature sets. At the same time, text-as-data research shows that NLP and LLMs can extract economically meaningful signals from large corpora, but their role as dedicated, standalone forecasting models for realised volatility remains underexplored. This motivates four objectives that are central to both researchers and practitioners. First, we assess whether news alone, without lagged realised volatility or other market variables, can generate competitive realised volatility forecasts, and whether general-purpose language representations are sufficiently competitive for this task or specialised finance-domain language representations are required. In doing so, we emphasise word embeddings as the core representation that converts unstructured news into quantitative inputs by encoding semantic and contextual similarity in a low-dimensional vector space; such embeddings also underpin modern LLMs as their representation layer, making them a transparent and practically relevant starting point for state-of-the-art NLP models in forecasting. Second, we examine whether any forecasting gains from news are stable across time and market conditions or concentrated in particular regimes. Third, we identify which components of news content drive predictability, distinguishing whether improvements arise primarily from stock-related or general news. Fourth, we evaluate the incremental value of combining news-based signals with standard volatility-history benchmarks relative to using either set of information in isolation. We address these objectives by developing a modern yet operationally simple NLP model that maps news directly into volatility forecasts. We evaluate it both as a news-only forecaster and as a complement to standard RV models, and use explainability analysis to identify the phrases and themes most responsible for predictive performance. We also extend the analysis to assess the economic gains of the proposed models.
@@ 3 Financial Word Embedding
Although general-purpose word embeddings have demonstrated strong performance in broad language tasks, it is important to consider the potential benefits of developing specialised word embeddings and to conduct a comparative evaluation with general-purpose alternatives at the language level. Financial language exhibits distinctive usage patterns, institutional references, and contextual relationships that are often absent from general text corpora. As a result, many tokens may carry meanings, associations, or economic implications that differ substantially from their usage in non-financial contexts. Using general-purpose word embeddings without prior validation may therefore introduce measurement error into text-based variables and weaken their economic interpretability. This consideration motivates the development and evaluation of specialised word embeddings tailored to financial language before employing them in downstream financial applications. Section 3.1 describes the steps from data collection to the training of the financial word embedding, together with the associated design considerations, and Section 3.2 subsequently presents the evaluation of both general-purpose and financial word embeddings using representation analysis and multiple language-level benchmarks.
@@ 3.1 Design
Turning to corpus size, the Google Word2Vec, WikiNews, and FinText corpora contain approximately 100 billion, 16 billion, and 4.32 billion words, respectively, implying that FinText relies on a substantially smaller corpus than these general-purpose word embeddings. Nevertheless, the FinText dataset expands substantially over time, with the number of monthly samples and the total word count exhibiting clear long-run growth over the 15-year period covered by the news archive, albeit with noticeable time variation, as illustrated in Figure 1. Following all pre-processing steps, the resulting FinText corpus comprises 2,733,035 unique tokens.
@@
3.2 Language-Level Evaluation
@@
Table 1 compares four versions of FinText against general-purpose word embeddings based on the Google Analogy benchmark. The Google Analogy benchmark is a widely used dataset for evaluating the quality of word embeddings. Each section in the Google Analogy benchmark contains a set of analogies. For example, under the ‘capital-common-countries’ section, the model is evaluated on analogies such as ‘London to England is like Paris to ?’. From Table 1, it is apparent that, except for ‘currency’ and ‘gram1-adjective-to-adverb’, WikiNews achieves the highest predictive accuracy. The overall evaluation score supports this observation. For this general-purpose benchmark, FinText is outperformed by Google under Word2Vec and by WikiNews under FastText. The individual and overall scores for FinText indicate that the skip-gram model performs better than CBOW.
@@ 4 Realised Volatility Forecasting
A substantial body of literature establishes that financial market volatility is fundamentally driven by the arrival and interpretation of information. Seminal contributions such as Engle and Ng (1993) and Andersen and Bollerslev (1998), together with subsequent intraday evidence, show that information arrival and return innovations are closely linked to volatility clustering and sudden volatility episodes (Gallo and Pacini 2000, Kalev et al. 2004). More broadly, this line of research emphasises that volatility dynamics reflect heterogeneous information flows operating across different horizons, which in turn generate the persistent behaviour commonly observed in financial markets (Andersen et al. 2007b). However, most empirical volatility forecasting models incorporate this information only indirectly, relying on past volatility measures as reduced-form summaries of how markets have responded to information arrivals. Although this approach has proven empirically successful, it abstracts from the underlying informational content and treats volatility persistence primarily as a time-series phenomenon. Recent evidence also shows that sentiment and attention measures from news and other sources provide additional predictive power for realised volatility beyond the HAR-family of models, although the gains are economically modest (Audrino et al. 2020). This motivates testing textual news as a direct proxy for information arrival.
@@ 5.2 Forecasting Performance of Ensemble Models
Although Section 5.1 shows that NLP models deliver promising results and that news content contains meaningful information for predicting RV across different market regimes, these models underperform the HAR-family of models in terms of standard forecasting metrics, particularly with respect to the magnitude of performance improvements. Combining forecasts from multiple models provides diversification gains and can improve predictive performance when individual forecasts rely on heterogeneous information sets or are subject to model uncertainty (Bates and Granger 1969, Timmermann 2006, Becker and Clements 2008). This naturally raises the question of whether, rather than serving as a replacement for the HAR-family of models, news-based NLP forecasts can provide complementary information. To this end, we define a straightforward ensemble model that combines forecasts from the CHAR model, which is the best-performing specification within the HAR-family in Section 5, with forecasts generated by the NLP models. For each day, we compute the average of the two forecasts. This approach integrates the persistent dynamics captured by historical RV with short term information conveyed by recent news, potentially enhancing the robustness of daily RV forecasts.
@@
Taken together, the ensemble results clarify and strengthen the conclusions drawn from the standalone NLP analysis in Section 5.1. While the NLP models typically underperformed the best HAR-family specifications in terms of loss ratios, the ensemble approach successfully converts the informational content embedded in news text into statistically meaningful forecasting gains. This complementarity is strongest for stock-related news, particularly in high volatility days, consistent with the earlier finding that stock-related news is most informative during periods of market stress. For general hot news, the ensemble results mirror the regime dependence identified previously, with benefits concentrated on normal volatility days. Across both news types, the ensemble evidence reinforces the earlier conclusion that performance improvements are driven primarily by the combination of heterogeneous information sources rather than by increased model complexity. This interpretation is consistent with the forecast combination literature, where Becker and Clements (2008) show in the context of volatility forecasting that combining forecasts can improve predictive accuracy relative to individual models. More importantly, these results indicate that news-based signals provide incremental predictive information for RV rather than fully substituting for volatility-history dynamics. Finally, similar to standalone NLP models in Section 5.1, specialised word embeddings perform better for stock-related news, while general-purpose word embeddings perform better for general hot news.
@@ 5.4 Explainable AI (XAI)
In LABEL:RV_movers_stock_related, we grouped the stock-related volatility drivers into ‘Analyst opinion’, ‘Event’, ‘Verb’, ‘Market’, ‘Abbreviation’, ‘Country/Company’, ‘Announcement’, ‘Numeric’, ‘Calendar’, ‘Insider’, and ‘Mixed’. [^27] It is important to note that this classification is based on our judgement. However, we believe that reassigning some of the n-grams between groups would not alter the fundamental informational content of these groups. The key findings can be summarised as follows: (i) ‘Analyst opinion’ and ‘Event’ contain the majority of volatility driver n-grams. This outcome is unsurprising, as it clearly demonstrates the importance of analyst opinions concerning a company’s earnings calls and financial reports. Among others, popular n-grams in this group include registers, announces, files, raises, and surrenders. (ii) ‘Market’ includes market-related n-grams such as stocks to buy, premarket, and stock market opens. (iii) China is the only country in the volatility drivers list for stock-related news, underscoring the relevance of news related to this country. The remaining classes contain fewer n-grams and exhibit less commonality; however, it is evident that these n-grams in stock-related news convey varying levels of information about stocks, the market, and the economy.
@@
LABEL:RV_movers_general_hot presents the volatility drivers in general hot news, grouped into ‘Person’, ‘Place’, ‘Legal entity’, ‘Level’, ‘Verb’, ‘Index’, ‘Data’, ‘Numeric’, and ‘Mixed’, and is revealing in several ways. The ‘Person’ category is dominated by prominent political and policy figures, including U.S. presidents Donald Trump, Barack Obama, and Joe Biden, New York City mayor Bill de Blasio, senior political actors such as Mark Meadows, Mitch McConnell, Mike Pompeo, and Nancy Pelosi, as well as key Federal Reserve officials, including Chair Jerome Powell, Janet Yellen, and Federal Reserve Bank presidents James Bullard of St. Louis, Loretta Mester of Cleveland, John C. Williams of New York, Neel Kashkari of Minneapolis, and Raphael Bostic of Atlanta, underscoring the central role of political and monetary policy communication in driving market volatility. International political figures such as Dominic Cummings in the United Kingdom and Kim Jong-un of North Korea also appear as important volatility drivers, while the presence of Jim Cramer, a prominent financial media commentator, is notable but unsurprising given the influence of news coverage on market sentiment. In contrast, the appearance of William G. Kaelin, a Nobel Prize-winning physician-scientist, and Reinhard Genzel, an astrophysicist, does not align with the economic and financial focus of the remaining n-grams in this category, illustrating that XAI methods are not error-freeand that their outputs must be interpreted with caution.
@@
Further analysis of LABEL:RV_movers_general_hot highlights the significance of places in our analysis and underscores the role of a group of countries, including China, as drivers of volatility. The term ‘Legal entity’ encompasses a variety of major offices, departments, commissions, and companies. The presence of the health organization and CDC (centres for disease control and prevention) is likely attributable to the COVID-19 pandemic. [^28] As expected, the word COVID-19 does not appear, as it was not in the corpus before 2015. The next significant class, with a high number of n-grams, is labelled ‘Level’, encompassing a range of levels and changes, percentages, currency values, and specific terms like below, above, fall, and under, all referencing certain quantitative expectations. Moving to the subsequent groups, similar to stock-related news, the ‘Verb’ and ‘Numeric’ groups in general hot news underscore the relevance of these particular verbs and numbers as volatility drivers. The ‘Mixed’ group includes a variety of n-grams, such as SPAC [^29] Special-purpose acquisition company., Payroll-tax cut, Airstrike, Shutdown, Coalitions, Crisis, Trade Speech, Attorney General, and Hearing. Finally, the ‘Index’ and ‘Data’ groups underscore the importance of changes in the S&P500 index and various financial and economic indicators such as inflation, GDP, and deficit as key n-grams. A detailed analysis of all n-grams is beyond the scope of this study; however, we believe that most of these n-grams align with expectations.